# Tree-Wasserstein Distance for High Dimensional Data with a Latent Feature Hierarchy

### v1

@@ 4.4 Binary Tree Construction
For sufficiently large $K_{c}$ and $m$, the decoded tree metric $d_{B}$ in Alg. 1 is bilipschitz equivalent to the underlying tree metric $d_{T}$, where $T$ is the ground latent tree assumed from Sec. 4.1 and $d_{T}$ is the hidden tree metric between features defined as the length of the shortest path on $T$. Specifically, the bilipschitz constants can be determined by $\frac{1}{\sqrt{2}}d_{B}(j_{1},j_{2})\leq d_{T}(j_{1},j_{2})\leq d_{B}(j_{1},j_{2})$ for all features $j_{1},j_{2}\in[m]$.
The proof of Thm. 4.1 is in App. C. It is important to recognize that the validity of Thm. 4.1 relies on the construction of the diffusion operator $\mathbf{P}$ from the observations $\mathbf{X}$, which is tightly connected to the heat kernel and the Laplacian of the underlying hierarchy $T$ (see Sec. 3). While $T$ is hidden, the diffusion operator $\mathbf{P}$ is used in our method to construct another tree, $B$, such that, the tree metrics of the data-driven tree $d_{B}$ and the hidden tree $d_{T}$ are equivalent. This result is stated in Thm. 4.1. In addition, our $\mathsf{HD}$ LCA for decoded tree $B$ is a continuous LCA in the product manifold of $(m+1)$-dimensional Poincaré half-spaces that recovers hidden tree relations. It is closely related to the hyperbolic LCA in Chami et al. [2020], where a continuous LCA in a single two-dimensional Poincaré disk was proposed for hierarchical clustering, and a tree was constructed by minimizing the Dasgupta’s cost [Dasgupta, 2016]. However, their method is not applicable to approximate $\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},T)$ since it does not recover the latent feature hierarchy, as we empirically show in Sec. 5.
@@ 4.5 Tree-Wasserstein Distance for High-Dimensional Data with a Latent Feature Hierarchy
where $\alpha_{v}=d_{B}(v,v^{\prime})$, and node $v^{\prime}$ is the parent of node $v$.
@@
For sufficiently large $K_{c}$ and $m$, the TWD defined in Eq. (3) is bilipschitz equivalent to the TWD with the ground latent tree $T$. Specifically, the bilipschitz constants can be determined by $\frac{1}{\sqrt{2}}\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},B)\leq\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},T)\leq\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},B)$ for all samples $i,i^{\prime}\in[n]$.
@@ 5 Experimental Results
We apply the proposed method to a synthetic example and to word-document and single-cell gene expression datasets. Additional implementation details are in App. D, including the hyperparameters and scaling to large datasets. Additional experiments, including runtime analysis, ablation studies, the impact of curvature, and robustness analysis under Gaussian noise, are in App. E.
@@ 5.1 Synthetic Word-Document Data Example
We compare the proposed TWD with existing TWDs, including (i) Quadtree [Indyk and Thaper, 2003] that constructs a random tree by a recursive division of hypercubes, (ii) Flowtree [Backurs et al., 2020] that computes the optimal flow of Quadtree, (iii) tree-sliced Wasserstein distance (TSWD) [Le et al., 2019] that computes an average of TWDs by random sampling trees, where the numbers of sampling are set to 1, 5, and 10, (iv) UltraTree [Chen et al., 2024] that constructs an ultrametric tree [Johnson, 1967] by minimizing OT regression cost, (v) weighted cluster TWD (WCTWD) [Yamada et al., 2022], (vi) weighted Quadtree TWD (WQTWD) [Yamada et al., 2022], and their sliced variants (vii) SWCTWD and (viii) SWQTWD [Yamada et al., 2022]. We also compare to TWDs where the trees are constructed by (i) MST [Prim, 1957], (ii) Tree Representation (TR) [Sonthalia and Gilbert, 2020] through a divide-and-conquer approach, (iii) gradient-based hierarchical clustering (HC) in hyperbolic space (gHHC) [Monath et al., 2019], (iv) gradient-based Ultrametric Fitting (UltraFit) [Chierchia and Perret, 2019], and (v) HC by hyperbolic Dasgupta’s cost (HHC) [Chami et al., 2020]. We refer to App. B for an extended description of these baselines.
@@ 5.2 Real-World Word-Document Data Analysis
Tab. LABEL:tab:classification presents the document classification accuracy. The results of Quadtree, Flowtree, TSWD, and WMD are taken from Huang et al. [2016], Takezawa et al. [2021]. Our method outperforms all the competing baselines on BBCSPORT, TWITTER, and AMAZON datasets, including the pre-trained WMD. On CLASSIC dataset, our proposed TWD yields the second-best classification accuracy and is competitive with WMD.
Table 1: Document and cell classification accuracy (the highest in bold and the second highest underlined). Results marked with are taken from Huang et al. [2016], Takezawa et al. [2021].
@@ Appendix B Additional Related Work
Wasserstein Distance on Hyperbolic Manifold. Wasserstein metric learning in hyperbolic spaces builds upon adapting the Wasserstein distance [Villani, 2009] to the structure of hyperbolic geometry [Lee, 2006]. This adaptation involves redefining the cost function in the optimization problem to incorporate the geodesic distance within the hyperbolic manifold. The resulting Wasserstein metric maintains its intuitive interpretation as a measure of the “effort” required to transform one distribution into another while under the manifold constraints of hyperbolic structure. It has proven valuable in various fields, for example, in shape analysis [Shi et al., 2016] and hierarchy matching [Alvarez-Melis et al., 2020, Hoyos-Idrobo, 2020]. Recently, the work in Bonet et al. [2023] proposed extending the sliced-Wasserstein distance [Rabin et al., 2012] to hyperbolic representation, enabling efficient computation of the Wasserstein distance. Our work, however, diverges fundamentally from these approaches. We focus on observational data that inherently possess a hierarchy rather than probability distributions defined on hyperbolic spaces. Additionally, our focus is on the TWD, aiming for efficient computation while addressing the specific challenges posed by the hidden feature hierarchy.
@@
Existing TWDs. The TWD is a popular method to reduce the computational complexity associated with calculating the Wasserstein distance. This approach approximates the original metric using a tree metric, as described in several studies [Indyk and Thaper, 2003, Le et al., 2019, Evans and Matsen, 2012, Leeb, 2018, Takezawa et al., 2021, Yamada et al., 2022]. The resulting methods provide a coarser approximation of OT in a Euclidean space, facilitating more efficient computations. One notable method is the Quadtree [Indyk and Thaper, 2003], which creates a random tree by recursively dividing hypercubes. This tree is then used to compute the TWD. Flowtree [Backurs et al., 2020] modifies the Quadtree method by focusing on the optimal flow and evaluating the cost of this flow within the ground metric. The Tree-Sliced Wasserstein Distance (TSWD) [Le et al., 2019] represents another TWD technique. It computes the average of TWDs by randomly sampling trees. The number of sampling trees is a hyperparameter, commonly set to one, five, or ten, resulting in the variants TSWD-1, TSWD-5, and TSWD-10. Further advancements are presented in Yamada et al. [2022], which introduce WQTWD and WCTWD. These methods employ a Quadtree or clustering-based tree to approximate the Wasserstein distances by optimizing the weights on the tree. Lastly, UltraTree [Chen et al., 2024] constructs an ultrametric tree [Johnson, 1967] by minimizing the OT regression cost, aiming to approximate the Wasserstein distance within the original metric space. Our approach, however, departs significantly from these existing methods. While we also aim to accelerate the computation of the Wasserstein distance, we leverage the tree structure to introduce and recover the latent feature hierarchy. Specifically, our novel hyperbolic tree decoding algorithm is designed to recover hidden feature hierarchies. Our TWD approximates the Wasserstein distance incorporating the latent feature hierarchy, rather than relying on the original metric in Euclidean space.
@@ Appendix C Theoretical Analysis and Proofs
For sufficiently large $K_{c}$ and $m$, the decoded tree metric $d_{B}$ in Alg. 1 is bilipschitz equivalent to the underlying tree metric $d_{T}$, where $T$ is the ground latent tree assumed from Sec. 4.1 and $d_{T}$ is the hidden tree metric between features defined as the length of the shortest path on $T$. Specifically, the bilipschitz constants can be determined by $\frac{1}{\sqrt{2}}d_{B}(j_{1},j_{2})\leq d_{T}(j_{1},j_{2})\leq d_{B}(j_{1},j_{2})$ for all features $j_{1},j_{2}\in[m]$.
@@
Therefore, the tree distance between $j_{1}$ and $j_{3}$ is given by $d_{B}(j_{1},j_{3})=d_{\mathcal{M}}(\mathbf{z}_{j_{1}},\mathbf{z}_{j_{1}}\vee\mathbf{z}_{j_{3}})+d_{\mathcal{M}}(\mathbf{z}_{j_{1}}\vee\mathbf{z}_{j_{3}},\mathbf{z}_{j_{2}}\vee\mathbf{z}_{j_{3}})+d_{\mathcal{M}}(\mathbf{z}_{j_{2}}\vee\mathbf{z}_{j_{3}},\mathbf{z}_{j_{3}})$. By triangle inequality, we have $d_{B}(j_{1},j_{3})\leq d_{\mathcal{M}}(\mathbf{z}_{j_{1}},\mathbf{z}_{j_{3}})$. We denote $d_{\mathcal{M}}(j_{1},j_{2})=d_{\mathcal{M}}(\mathbf{z}_{j_{1}},\mathbf{z}_{j_{3}})$ for simplicity. By Prop. 4.2 and Prop. 4.4, we have $c\cdot d_{\mathcal{M}}({j_{1}},{j_{3}})\leq d_{B}(j_{1},j_{3})$. Moreover, since the HD-LCA relation $[j_{2}\vee j_{3}\sim j_{1}]_{\mathcal{M}}$ holds, the tree distance between $j_{2},j_{3}\in[m]$ on $B$ can be directly obtained by $d_{B}(j_{2},j_{3})=d_{\mathcal{M}}(j_{2},j_{3})$. As only one of the HD-LCA relations $[j_{2}\vee j_{3}\sim j_{1}]_{\mathcal{M}}$, $[j_{1}\vee j_{3}\sim j_{2}]_{\mathcal{M}}$, or $[j_{1}\vee j_{2}\sim j_{3}]_{\mathcal{M}}$ holds in $B$, the proof can be applied to any three points $j_{1},j_{2},j_{3}\in[m]$ of which HD-LCA relation holds. Using Monte-Carlo integration, the convergence is with probability one as $K_{c}\rightarrow\infty$. For a finite set, we have a high probability bound on the convergence. Therefore, $d_{B}\simeq d_{\mathcal{M}}\simeq d_{T}$. Specifically, by taking $\alpha\rightarrow\frac{1}{2}$ and $\beta=2$ in Section 3 in Leeb and Coifman [2016] and Appendix C in Lin et al. [2023], the bilipschitz constants $c$ and $C$ can be determined by $\frac{1}{\sqrt{2}}$ and $1$ in $d_{B}\simeq d_{T}$, respectively. ∎
@@
For $\alpha\rightarrow\frac{1}{2}$ and sufficiently large $K_{c}$, we have $d_{\mathcal{M}}\simeq d_{T}$.
@@
For sufficiently large $K_{c}$ and $m$, the TWD defined in Eq. (3) is bilipschitz equivalent to the TWD with the ground latent tree $T$. Specifically, the bilipschitz constants can be determined by $\frac{1}{\sqrt{2}}\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},B)\leq\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},T)\leq\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},B)$ for all samples $i,i^{\prime}\in[n]$.
@@
Note that based on Thm. 4.1, the decoded tree metric $d_{B}(\cdot,\cdot)$ and the underlying tree metric $d_{T}(\cdot,\cdot)$ are bilipschitz equivalent. Additionally, the TWD using the tree $T$ is the Wasserstein distance with ground pairwise distance $d_{T}$ [Evans and Matsen, 2012], i.e., $\mathtt{TW}(\mu_{1},\mu_{2},T)=\mathtt{OT}(\mu_{1},\mu_{2},d_{T})$. Therefore, for any two discrete histograms $\mathbf{x}_{i}$ and $\mathbf{x}_{i^{\prime}}$ supported on $T$, we have $\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},B)\simeq\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},T)$, and the bilipschitz constants are $\frac{1}{\sqrt{2}}$ and $1$, respectively. ∎

### v2

@@ 4.4 Binary Tree Construction
For sufficiently large $K_{c}$ and $m$, there exist some constants $C_{1},C_{2}>0$ such that $C_{1}d_{T}(j_{1},j_{2})\leq d_{B}(j_{1},j_{2})\leq C_{2}d_{T}(j_{1},j_{2})$ for $j_{1},j_{2}\in[m]$, where $d_{B}$ is the decoded tree metric $d_{B}$ in Alg. 1, $T$ is the ground truth latent tree assumed from Sec. 4.1, and $d_{T}$ is the hidden tree metric between features defined as the length of the shortest path on $T$ (i.e., the geodesic distances).
The proof of Thm. 4.1 is in App. C. The validity of Thm. 4.1 relies on the assumption that the local Gaussian affinities between features allow us to reveal the hidden hierarchy. Formally, in the proof of Thm. 4.1, we assume that the discrete diffusion operator $\mathbf{P}$ constructed based on the local Gaussian affinities between features in the ambient space $\mathbb{R}^{n}$, in the limit of $m\rightarrow\infty$ and $\epsilon\rightarrow 0$, converges pointwise to the heat kernel of the underlying hierarchical metric space $(\mathcal{H},d_{\mathcal{H}})$. While $T$ is hidden, the diffusion operator $\mathbf{P}$ is accessible and used to construct another tree, $B$, such that, the tree distances of the inferred tree $d_{B}$ and the hidden tree $d_{T}$ are equivalent. This result is stated in Thm. 4.1. In addition, our HD-LCA is a continuous LCA in the product manifold of $(m+1)$-dimensional Poincaré half-spaces that recovers hidden tree relations. It is closely related to the hyperbolic LCA in Chami et al. [2020], where a continuous LCA in a single 2D Poincaré disk was proposed for hierarchical clustering, and a tree was constructed by minimizing the Dasgupta’s cost [Dasgupta, 2016]. However, their method does not recover the latent feature hierarchy, as shown in Sec. 5.
@@ 4.5 Tree-Wasserstein Distance for High-Dimensional Data with a Latent Feature Hierarchy
where $B$ is the decoded tree in Alg. 1, $\alpha_{v}=d_{B}(v,v^{\prime})$, and node $v^{\prime}$ is the parent of node $v$.
@@
For sufficiently large $K_{c}$ and $m$, there exist some constant $\widetilde{C}_{1},\widetilde{C}_{2}>0$ such that $\widetilde{C}_{1}\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},T)\leq\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},B)\leq\widetilde{C}_{2}\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},T)$ for all samples $i,i^{\prime}\in[n]$.
@@ 5 Experimental Results
We apply the proposed method to a synthetic example and to word-document and single-cell RNA-sequencing (scRNA-seq) data. We use Word2vec [Mikolov et al., 2013] as word embedding vectors for the word-document data and Gene2vec [Du et al., 2019] as gene embedding vectors in scRNA-seq experiments. The implementation details are in App. D. Additional experiments are in App. E.
@@ 5.1 Synthetic Word-Document Data Example
We compare the proposed TWD with existing TWDs, including (i) Quadtree [Indyk and Thaper, 2003] that constructs a random tree by a recursive division of hypercubes, (ii) Flowtree [Backurs et al., 2020] that computes the optimal flow of Quadtree, (iii) tree-sliced Wasserstein distance (TSWD) [Le et al., 2019] that computes an average of TWDs by random sampling trees, where the numbers of sampling are set to 1, 5, and 10, (iv) UltraTree [Chen et al., 2024] that constructs an ultrametric tree [Johnson, 1967] by minimizing OT regression cost, (v) weighted cluster TWD (WCTWD) [Yamada et al., 2022], (vi) weighted Quadtree TWD (WQTWD) [Yamada et al., 2022], and their sliced variants (vii) SWCTWD and (viii) SWQTWD [Yamada et al., 2022]. We also compare to TWDs where the trees are constructed by (i) MST [Prim, 1957], (ii) Tree Representation (TR) [Sonthalia and Gilbert, 2020] through a divide-and-conquer approach, (iii) gradient-based hierarchical clustering (HC) in hyperbolic space (gHHC) [Monath et al., 2019], (iv) gradient-based Ultrametric Fitting (UltraFit) [Chierchia and Perret, 2019], and (v) HC by hyperbolic Dasgupta’s cost (HHC) [Chami et al., 2020]. We refer to App. B for a detailed description of baselines and highlight the novel aspects of our TWD.
@@ 5.2 Real-World Word-Document Data Analysis
Tab. LABEL:tab:classification presents the document classification accuracy. The results of Quadtree, Flowtree, TSWD, and WMD are taken from Huang et al. [2016], Takezawa et al. [2021]. Our method outperforms all the competing baselines on BBCSPORT, TWITTER, and AMAZON datasets, including the pre-trained WMD that is specifically designed for word-document data. On CLASSIC dataset, our proposed TWD yields the second-best classification accuracy and is competitive with WMD.
Table 1: Document and cell classification accuracy (the highest in bold and the second highest underlined). Results marked with are taken from Huang et al. [2016], Takezawa et al. [2021]. The empirical comparisons to TWD methods demonstrate the advantages of our proposed approach, which effectively incorporates the latent feature hierarchy, leading to improved performance.
@@ A.2 Diffusion Geometry
Note that the density-normalized affinity matrix was considered to mitigate the effect of non-uniform data sampling, which was proposed and shown theoretically in Coifman and Lafon [2006]. In Section 3.1 in Coifman and Lafon [2006], the construction of a family of diffusion operators is introduced with a parameter $\alpha\in\mathbb{R}$ and the density $q$. In Section 3.4 in Coifman and Lafon [2006], it is shown that when $\alpha=1$ (in our case $\widehat{\mathbf{Q}}=\mathbf{D}^{-1}\mathbf{Q}\mathbf{D}^{-1}$), the diffusion operator approximates the heat kernel, and this approximation is independent of the density $q$. This technique has been further explored and utilized in [Mishne et al., 2019, Katz et al., 2020, Shnitzer et al., 2022, 2024].
@@ A.4 Preliminaries on Trees
Definition A.8 (LCA Relation [Wang and Wang, 2018].).
Given a tree $T$, we say that the LCA relation $[j_{2}\vee j_{3}\sim j_{1}]_{T}$ holds in $T$, if the LCA of $j_{2}$ and $j_{3}$, denoted as $j_{2}\vee j_{3}$, is a proper descendant of the LCA of $j_{2}$, $j_{3}$ and $j_{1}$, denoted as $j_{2}\vee j_{3}\vee j_{1}$.
Remark.
While very small or specific types (e.g., star-shaped) of trees may be isometrically embedded in Euclidean space, general trees or hierarchical structures cannot, and embedding them into Euclidean spaces results in larger distortion. For larger or more complex trees, hyperbolic space often serves as a more natural embedding space, preserving their hierarchical structure and distances.
@@ Appendix B Additional Related Work
Wasserstein Distance on Hyperbolic Manifold. Wasserstein distance learning in hyperbolic spaces builds upon adapting the Wasserstein distance [Villani, 2009] to the structure of hyperbolic geometry [Lee, 2006]. This adaptation involves redefining the cost function in the optimization problem to incorporate the geodesic distance within the hyperbolic manifold. The resulting Wasserstein metric maintains its intuitive interpretation as a measure of the “effort” required to transform one distribution into another while under the manifold constraints of hyperbolic structure. It has proven valuable in various fields, for example, in shape analysis [Shi et al., 2016] and hierarchy matching [Alvarez-Melis et al., 2020, Hoyos-Idrobo, 2020]. Recently, the work in Bonet et al. [2023] proposed extending the sliced-Wasserstein distance [Rabin et al., 2012] to hyperbolic representation, enabling efficient computation of the Wasserstein distance. Our work, however, diverges fundamentally from these approaches. We focus on observational data that inherently possess a hierarchy rather than probability distributions defined on hyperbolic spaces. Additionally, our focus is on the TWD, aiming for efficient computation while addressing the specific challenges posed by the hidden feature hierarchy.
@@
Existing TWDs. The TWD is a popular method to reduce the computational complexity associated with calculating the Wasserstein distance on the Euclidean metric. This approach approximates the original Euclidean metric using a tree metric, as described in several studies [Indyk and Thaper, 2003, Le et al., 2019, Evans and Matsen, 2012, Leeb, 2018, Takezawa et al., 2021, Yamada et al., 2022]. The resulting methods provide a coarser approximation of OT in an Euclidean space, facilitating more efficient computations from cubic complexity to linear complexity. One notable method is the Quadtree [Indyk and Thaper, 2003], which creates a random tree by recursively dividing hypercubes. This tree is then used to compute the TWD. Flowtree [Backurs et al., 2020] modifies the Quadtree method by focusing on the optimal flow and evaluating the cost of this flow within the ground metric. The Tree-Sliced Wasserstein Distance (TSWD) [Le et al., 2019] represents another TWD technique. It computes the average of TWDs by randomly sampling trees. The number of sampling trees is a hyperparameter, commonly set to one, five, or ten, resulting in the variants TSWD-1, TSWD-5, and TSWD-10. Further advancements are presented in Yamada et al. [2022], which introduce WQTWD and WCTWD. These methods employ a Quadtree or clustering-based tree to approximate the Wasserstein distances by optimizing the weights on the tree. Lastly, UltraTree [Chen et al., 2024] constructs an ultrametric tree [Johnson, 1967] by minimizing the OT regression cost, aiming to approximate the Wasserstein distance within the original metric space.
Our approach, however, departs significantly from these existing methods. While we also aim to accelerate the computation of the Wasserstein distance using a tree, we leverage the tree structure to introduce and recover the latent feature hierarchy. Specifically, our novel hyperbolic tree decoding algorithm is designed to recover hidden feature hierarchies. Our TWD approximates the Wasserstein distance incorporating the latent feature hierarchy, rather than relying on the original metric in Euclidean space.
@@ Appendix C Theoretical Analysis and Proofs
For sufficiently large $K_{c}$ and $m$, there exist some constants $C_{1},C_{2}>0$ such that $C_{1}d_{T}(j_{1},j_{2})\leq d_{B}(j_{1},j_{2})\leq C_{2}d_{T}(j_{1},j_{2})$ for $j_{1},j_{2}\in[m]$, where $d_{B}$ is the decoded tree metric $d_{B}$ in Alg. 1, $T$ is the ground truth latent tree assumed from Sec. 4.1, and $d_{T}$ is the hidden tree metric between features defined as the length of the shortest path on $T$ (i.e., the geodesic distances).
@@
Therefore, the tree distance between $j_{1}$ and $j_{3}$ is given by $d_{B}(j_{1},j_{3})=d_{\mathcal{M}}(\mathbf{z}_{j_{1}},\mathbf{z}_{j_{1}}\vee\mathbf{z}_{j_{3}})+d_{\mathcal{M}}(\mathbf{z}_{j_{1}}\vee\mathbf{z}_{j_{3}},\mathbf{z}_{j_{2}}\vee\mathbf{z}_{j_{3}})+d_{\mathcal{M}}(\mathbf{z}_{j_{2}}\vee\mathbf{z}_{j_{3}},\mathbf{z}_{j_{3}})$. By triangle inequality, we have $d_{B}(j_{1},j_{3})\leq d_{\mathcal{M}}(\mathbf{z}_{j_{1}},\mathbf{z}_{j_{3}})$. We denote $d_{\mathcal{M}}(j_{1},j_{2})=d_{\mathcal{M}}(\mathbf{z}_{j_{1}},\mathbf{z}_{j_{3}})$ for simplicity. By Prop. 4.2 and Prop. 4.4, we have $c\cdot d_{\mathcal{M}}({j_{1}},{j_{3}})\leq d_{B}(j_{1},j_{3})$. Moreover, since the HD-LCA relation $[j_{2}\vee j_{3}\sim j_{1}]_{\mathcal{M}}$ holds, the tree distance between $j_{2},j_{3}\in[m]$ on $B$ can be directly obtained by $d_{B}(j_{2},j_{3})=d_{\mathcal{M}}(j_{2},j_{3})$. As only one of the HD-LCA relations $[j_{2}\vee j_{3}\sim j_{1}]_{\mathcal{M}}$, $[j_{1}\vee j_{3}\sim j_{2}]_{\mathcal{M}}$, or $[j_{1}\vee j_{2}\sim j_{3}]_{\mathcal{M}}$ holds in $B$, the proof can be applied to any three points $j_{1},j_{2},j_{3}\in[m]$ of which HD-LCA relation holds. Using Monte-Carlo integration, the convergence is with probability one as $K_{c}\rightarrow\infty$. For a finite set, we have a high probability bound on the convergence. Therefore, $d_{B}\simeq d_{\mathcal{M}}\simeq d_{T}$.
@@
For $\alpha\rightarrow\frac{1}{2}$ and sufficiently large $K_{c}$, there exists some constants $\widetilde{C}_{1},\widetilde{C}_{2}>0$ such that $\widetilde{C}_{1}d_{T}^{2\alpha}\leq d_{\mathcal{M}}\leq\widetilde{C}_{2}d_{T}^{2\alpha}$.
@@
For sufficiently large $K_{c}$ and $m$, there exist some constant $\widetilde{C}_{1},\widetilde{C}_{2}>0$ such that $\widetilde{C}_{1}\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},T)\leq\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},B)\leq\widetilde{C}_{2}\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},T)$ for all samples $i,i^{\prime}\in[n]$.
@@
Note that based on Thm. 4.1, the decoded tree metric $d_{B}(\cdot,\cdot)$ and the underlying tree metric $d_{T}(\cdot,\cdot)$ are bilipschitz equivalent. Additionally, the TWD using the tree $T$ is the Wasserstein distance with ground pairwise distance $d_{T}$ [Evans and Matsen, 2012], i.e., $\mathtt{TW}(\mu_{1},\mu_{2},T)=\mathtt{OT}(\mu_{1},\mu_{2},d_{T})$. Therefore, for any two discrete histograms $\mathbf{x}_{i}$ and $\mathbf{x}_{i^{\prime}}$ supported on $T$, we have $\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},B)\simeq\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},T)$. ∎